Determine the area under the standard normal distribution curve between z=1.15 and z=-1.85

1
Expert's answer
2022-01-30T16:45:03-0500

"P(-1.85<Z<1.15) = P(Z<1.15) –P(Z<-1.85)\\
=0.8748-0.0322=0.8427\\"